OVERVIEW OF THE SPFIT PROGRAM SET

\underline{HERBERT M. PICKETT}, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California 91109.

The SPFIT program can fit spectral lines for molecules with up to 99 vibronic states and up to 10 spins with a variety of different operators. The SPCAT program can predict the intensities and positions of such a molecule with a variety of electric and magnetic dipole matrix elements. SPFIT can accommodate blended lines and line frequencies in units of MHz and cm^{−1}. SPCAT produces a line list in JPL catalog format and can also produce an energy listing that includes the estimated uncertainty and fractional perturbation. I will give a brief description will be given of some of the algorithms used internally and will address some frequently asked questions about how to use the program.